OpenStudy (anonymous):

Find the next three terms in the sequence 625, 250, 100, 40,

hartnn (hartnn):

did you notice that there is a common ratio here ? 625/250 =... ? 250/100 =..? 100/40=...? that would be your common ratio 'r' which equals ?

hartnn (hartnn):

when you find that 'r' you can use it to find any term! next term = 40/r= ...? (say 'x') 2nd next term = x/r =...? (say 'y') 3rd nest term = y/r = ....?